JAHN-TELLER ANALYSIS OF THE VIBRONIC STRUCTURE OF THE X̃^{2}E′′_{1} STATE IN THE CYCLOPEN-TADIENYL RADICAL

\underline{BRIAN E. APPLEGATE}, ANDREW J. BEZANT, AND TERRY A. MILLER, The Ohio State University, Dept. of Chemistry, Laser Spectroscopy Facility, 120 W. 18th Avenue, Columbus, Ohio 43210; TIMOTHY A.BARCKHOLTZ, JILA, National Institute of Standards and Technology and The Department of Chemistry and Biochemistry, University of Colorado Boulder, Colorado 80309.

While room temperature wavelength resolved emission spectra of the Ã^{2}A′′_{2}-X̃^{2}E′′_{1} transition have existed for nearly 20 years^{a}, the vibrational assignment of these spectra has remained elusive. The major difficulty with the vibrational analysis is attributable to complications arising from the dynamic Jahn-Teller effect in the X̃ state. Newly obtained jet-cooled laser excited wavelength resolved fluorescence emmision spectra, in conjunction with recent ^{b}calculations aimed at predicting the relevant Jahn-Teller constants have now made the complete analysis of the available spectral data possible. The transitions involving the Jahn-Teller active vibrations have been analyzed in terms of the three lowest energy harmonic vibrations of the appropriate symmetry (e′_{2}), assuming only linear Jahn-Teller interactions. Additional features of the spectrum may be described in terms of the fundamentals, overtones, and combination bands of the non-Jahn-Teller active vibrations as well as combinations involving the Jahn-Teller active modes.

aNelson H. H.; Pasternack L.; McDonald J. R., Chem. Phys, 1983, 74, 227.

bApplegate B. E.; Barckholtz T. A.;Miller T. A., to be published